Phineas is pushing a 11.13-kg box of tools with a force of 47 N directed at an angle of 83° downward from the horizontal. How fast is the cart going after the acceleration from rest if their backyard is 77 cm away from the garage?


1
Expert's answer
2022-03-17T09:30:58-0400

The horizontal component of applied force

"F_x=F\cos\theta\\
=47\:\rm N*cos83^\circ=5.73\: N"

The acceleration of the box

"a=\frac{F_x}{m}=\frac{5.73\:\rm N}{11.13\:\rm kg}=0.515\:\rm m/s^2"

The final velocity

"v=\sqrt{2ad}\\
=\sqrt{2*0.515\:\rm m/s^2*0.77\: m}=0.89\:\rm m/s"